OpenStudy (anonymous):

Situation: A hiker in africa discovers a skull that contains 43% of it's original c-14. k = .0001 find the age of the skull, round to the nearest year

OpenStudy (anonymous):

... already answered Please stop spamming :D

OpenStudy (anonymous):

no it's not already answered

OpenStudy (anonymous):

That's my bad, I did the work, but must not have posted. REALLY SORRY: \[f(t) = Ae^{-kt}\] \[f(t)/A = .43\] \[e^{-kt} = .43; -kt = \ln(.43); t = \ln(.43)/-k\] t = 8439.7007 = 8440 years

OpenStudy (anonymous):

thank you